If the measure of the three angles of a triangle are x, 2x-20, and 3x-10, then what it this type of trangle? Why? I don't understand how to do this problem could someone please help me?
Answer by edjones(7311)

there are 180 deg. in any triangle.
x+2x-20+3x-10=180
6x=180+30
6x=210
x=35 deg.
2x-20=50 deg.
3x-10=95 deg.
Ed